You will be asked to find the simple interest for an account.
You have $10,000 in a savings account that pays 2% simple interest per year. Find the
interest earned for the first year.
Find the balance on an account that accrues compound interest . Given
an investment, APR, compounding rate and length of investment find the balance after maturity.
You have $10,000 in a savings account that pays 2% APR compounded monthly. Find the
balance on the account after 5 years
Answer #1

1.
=Value*rate
=10000*2%=200

2.
=Value*(1+rate)
=100000*1.02=102000

3.
=Value*(1+r/m)^(m*t)
=10000*(1+2%/12)^(12*5)=11050.7892653082